OpenStudy (anonymous):

218 students went on a field trip. Five buses were filled and 28 students traveled in cars. How many students were in each bus? WRITE THE EQUATION

OpenStudy (anonymous):

5x+28 =218 solve.

OpenStudy (anonymous):

assuming each bus holds the same number of people.

OpenStudy (stormfire1):

The number of students per bus (n) = (218-28)/5 n = 38